OpenStudy (anonymous):

Solve the following equation in a word document. Be sure to show all work. You should have a minimum of 7 lines of work. 5(2x - 7) = 33 - 3(6x + 4)

OpenStudy (radar):

Clear the parenthesis getting 10x -35=33-18x-12 now add 35 to both sides getting: 10x=33+35-18x-12, now add 18x to both sides getting: 10x+18x=33+35-12, now combine similar terms, getting: 28x=56, now divide both sides by 28 getting: x=2

OpenStudy (anonymous):

\[5(2x-7)=33-3(6x+4)\]\[10x-35=33-18x-12\]\[10x+18x=33+35-12\]\[28x=56\]\[x=56/28\]\[x=2\]

OpenStudy (radar):

you could make 7 lines putting the equation on separate lines from the explanation/directions etc.